OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

Find slope. \[m = \frac{ y2 - y1 }{ x2 - x1 }\]\[m = \frac{ 2 - (-4) }{ -2 - 0 } = \frac{ 2 + 4 }{ -2 - 0 } = \frac{ 6 }{ -2 } = -3\]Slope is -3. You have the slope(-3) and you have a given point (0, -4). Point-Slope form. y - y1 = m(x - x1) m is slope Substitute the x- and y-values of your point, for x1 and y1. So you have:\[y - (-4) = -3(x - 0)\]or\[y + 4 = -3(x - 0)\] Change into Slope-Intercept form. y = mx + b \[y + 4 = -3(x - 0)\]Distribute. \[y + 4 = -3x - 0\]Solve for y. Subtract 4 from both sides.\[y = -3x - 4\]^---There is your equation.

OpenStudy (anonymous):

Hopefully this helped you understand the concept! (: